The children at Tiny Tots Day care plant a rectangular vegetable garden with a perimeter of 39 m. The length is twice the width. Find the dimensions of the garden

Found 2 solutions by stanbon, rvquartz:
Answer by stanbon(75887)

The children at Tiny Tots Day care plant a rectangular vegetable garden with a perimeter of 39 m. The length is twice the width. Find the dimensions of the garden.
---
Let the width be "x".
Then the length is "2x".
----
Equation:
Perimeter = 2(width + length)
39 = 2(x+2x)
39 = 2*3x
13 = 2x
x = 6.5m (width)
---
2x = 13m (length)
======================
Cheers,
Stan H.
---------------

Answer by rvquartz(19)

2 times the length + 2 times the width = 39 meters
Let L = 'length'
Let W = 'width'
we can now model the problem as:
2L + 2W = 39
from the problem statement we know that L = 2W so,
2(2W) + 2W = 39, which is the same as saying
4W + 2W = 39, which is the same as saying
6W = 39
W = 39/6
W = 6.5 meters and since L = 2W,
L = 13 meters